OpenStudy (anonymous):

Help for a medal c: Explain how to convert f(x) into the general, vertex form of the equation. Use complete sentences. You may use the f(x)=x^2-4

OpenStudy (usarmy3947):

im sorry i cant help but my friends can: @AllTheLonelyKillers @texaschic101 @Dobby1 @Compassionate @Goomblet @elliecosme

OpenStudy (anonymous):

Aw ok, thanks c:

OpenStudy (usarmy3947):

np : )

OpenStudy (compassionate):

Does this relate to quadratics?

OpenStudy (anonymous):

Um yes (:

OpenStudy (compassionate):

The general form of a quadratic is "y = ax2 + bx + c". For graphing, the leading coefficient "a" indicates how "fat" or how "skinny" the parabola will be. For | a | > 1 (such as a = 3 or a = –4), the parabola will be "skinny", because it grows more quickly (three times as fast or four times as fast, respectively, in the case of our sample values of a). For | a | < 1 (such as a = 1/3 or a = –1/4 ), the parabola will be "fat", because it grows more slowly (one-third as fast or one-fourth as fast, respectively, in the examples). Also, if a is negative, then the parabola is upside-down.

OpenStudy (anonymous):

wait idek tbh im so confused on this. the first question was this: To open these doors, you must speak three functions in standard form. One function, f(x), with two real rational solutions. f(x)=x2-4 One function, g(x), with two real irrational solutions. g(x)= x2-2 One function, h(x), with two complex solutions. h(x)= -3x2+11i and im not really sure if my answers or right or not.

OpenStudy (anonymous):

Ok yeah i get that part (: but how do i covert it? x

OpenStudy (anonymous):

*convert

OpenStudy (usarmy3947):

um @omgkelley i sent u a pm just let me know if u need further help with this : )

OpenStudy (anonymous):

a what?

OpenStudy (compassionate):

PM - Private Message

OpenStudy (anonymous):

Oh okay haha thanks !